Raw Cotton $28.00

Spinning/Weaving/Dyeing $12.00

Cut/Sew/Finishing $10.00

Material Transportation $ 3.00

Factory Fees $16.00

Inspection and Import Fees $14.00

Ocean Freight/Insurance $ 15.00

Warehousing $ 8.00

Packaging $15.00

Promotion $30.00

Customer Shipping $15.00

Given the cost per king-size sheet set above, and assuming the manufacturer has total fixed costs of $600,000 and estimates first year sales will be 50,000 sets Answer the following questions:

Determine the price to consumers if the company desires a 50 percent mark-upon cost. (15 marks)

 

1
Expert's answer
2021-05-31T17:01:02-0400

Answer

Cost per king sheet set will be 

$28 + $12 + $10 + $3 + $16 + $14 + $5 + $8 + $15 + $30 + $15 = $ 156 


Evaluated first year deals = 50,000 sets 


All out fixed expense = $600,000 

Normal fixed expense per extra large sheet set = 600,000/50,000 = $12

Absolute Cost per extra large sheet set = $156 + $12 = $168


Wanted edge on deals = 50 % 

Let us consider the deal cost to be $100x 

On the off chance that edge is 50% of the deals that implies edge = (50/100)*100x = 50x 

Along these lines, cost ought to be= $(100 - 50) = $60x 

Likewise, cost = $168


→ 168 = 60x 


→ x = 166/60 = 2.8


In this way, deal cost = $(100*2.8) = $280appx.


10-14 


Presently, Company offers first to the distributer who at that point offers it to retailer. Retailer at that point offers it to clients at a cost of $288


Since retailer gets the edge of 20%, So the edge of retailer will be = (20/100)*288 = $57.6


Henceforth the cost at which retailer gets the sheet set from distributer will be = $(288-57.6) = $230.4

Distributer sells the sheet set to the retailer for $230.4 and gets the edge of 10% 


→ edge of the distributer = $(10/100)*230.4 = $23.4


Thus the cost at which distributer gets the sheet set = $(230.4 - 23.4) = $207

In this way, cost at which organization currently offers to the distributer will be $207